OpenStudy (anonymous):

Just looking for someone to check if my answer is right. Calculate the derivative: y=sqrt2 +2/x at yprime(16) I had 1/2

OpenStudy (anonymous):

\[y=\frac{ \sqrt{x}+2 }{ x }\]

OpenStudy (anonymous):

\[y ^{'}(16)\]

jimthompson5910 (jim_thompson5910):

What did you get for y ' ?

OpenStudy (anonymous):

i think i got it wrong, because it's pretty wonky. \[\frac{ x \sqrt{x}-4\sqrt{x} }{ 2\sqrt{x} } \times \frac{ 1 }{ x ^{2} }\]

OpenStudy (anonymous):

i used the quotient rule - \[\frac{ x \times \frac{ 1 }{ 2\sqrt{x} }-\sqrt{x}+2(1) }{ x ^{2} }\]

jimthompson5910 (jim_thompson5910):

I like the product rule (it's easier to remember in my opinion) y = (sqrt(x) + 2)/x y = (sqrt(x) + 2)*x^(-1) y' = d/dx[(sqrt(x) + 2)]*x^(-1) + (sqrt(x) + 2)*d/dx[x^(-1)] y' = 1/(2*sqrt(x))*x^(-1) + (sqrt(x) + 2)*(-x^(-2)) You can simplify a number of ways, but you can jump to the part where you plug in x = 16 and evaluate to get... y'(16) = 1/(2*sqrt(16))*(16)^(-1) + (sqrt(16) + 2)*(-(16)^(-2)) y'(16) = -1/64 .... or .... y'(16) = -0.015625

OpenStudy (anonymous):

how can i use the product rule when it's division?

jimthompson5910 (jim_thompson5910):

the quotient x/y can be rewritten as a product x*y^(-1)

jimthompson5910 (jim_thompson5910):

since 1/y = y^(-1)

OpenStudy (anonymous):

so i change it into a product then i can just do the product rule, can i only do this when there is just one variable in the denominator?

jimthompson5910 (jim_thompson5910):

you can do it for any algebraic expression in the denominator A more complicated example \[\Large \frac{x^2+3x}{3x^3+9x^2+7x}\] is the same as \[\Large (x^2+3x)(3x^3+9x^2+7x)^{-1}\]

jimthompson5910 (jim_thompson5910):

You just write the numerator and denominator as a product and you write an exponent of -1 over the second part

OpenStudy (anonymous):

i see!

jimthompson5910 (jim_thompson5910):

of course, the quotient rule still works (I just forget it all the time and have to look it up constantly...so I rewrite the quotient into a product and use the product rule since it's easier to remember for me)